Determine whether or not ∗ gives a group structure on the set. If it is not a group, say which

axioms fail to hold.

Define ∗ on Z by a ∗ b = ab.


1
Expert's answer
2021-07-05T17:18:22-0400

Consider the operation "\ast" on "Z" defined by

"a\ast b=ab"

Require to determine whether or not "\ast" gives a group structure on the set "Z"

Let us verify the group axioms:

(1) Closure: "a,b\in Z\Rightarrow a*b\in Z"

Now "a,b\in Z\Rightarrow ab\in Z"

"\Rightarrow a* b\in Z"

So, "a,b\in Z\Rightarrow a*b\in Z"

Therefore, "Z" satisfies the closure axiom under the operation "\ast"

(2) Associativity: "a,b,c\in Z\Rightarrow a* (b* c)=(a* b)\star c"

Now "a,b,c\in Z\Rightarrow a* (b* c)=a* (bc)=a(bc)=abc"

And

"(a* b)* c=(ab)* c=(ab)c=abc"

So, "a,b,c\in Z\Rightarrow a* (b* c)=(a*b)* c"

Therefore, "Z" satisfies the associativity axiom under the operation "\ast"

(3) Identity: For each "a\in Z" there exists "e\in Z" such that "a*e=a=e*a"

Now "a*e=a\Rightarrow ae=a"

"\Rightarrow e=1\in Z"

And "a*1=a(1)=a=1(a)=1*a"

Therefore, "1" is the identity in "Z"

(4) Inverse: For each non zero "a\in Z" there exists "b\in Z" such that "a*b=e=b*a"

Let "a" be a non zero integer

Now "a*b=1\Rightarrow ab=1"

"\Rightarrow b=\frac{1}{a}\notin Z"

Therefore, inverse axiom fail to hold

Hence, "(Z,*)" is not a group.